what is 1 and 3/4 with the denominator of 12 siplafied
Answer by ptaylor(2198)

1 and 3/4=4/4+3/4=7/4
(7/4)/12 Multiply numerator and denomiator by (1/12). This will get rid of the complex fraction
(7/4)(1/12)/12(1/12)=
7/48/1=7/48----------------answer
Hope this helps---ptaylor